# supplemental

### v2

@@ Abstract
Figure 1: We introduce MaterialPicker, a DiT-based model that generates high-quality materials, conditioned on image crops and/or text prompts. Our model accurately captures textures and material properties even from photographs of distorted or partially obscured surfaces. We demonstrate MaterialPicker by extracting material properties (albedo, normal, height, roughness and metallicity, shown in a column next to the input crops) from smartphone-captured photos, then applying these materials in a 3D scene [^2] The 3D model in Fig. 2 was downloaded from Open3DLab [33], licensed under CC BY-NC-ND 4.0 [7]. We gratefully acknowledge Open3DLab for providing the model. For the purpose of visualization in our work, we made adjustments to the positions of certain objects, modified the material maps, and updated the lighting settings in Blender.for photo-realistic rendering results.
1 Introduction
@@
Traditional material acquisition often requires tens or hundreds of photo samples under known light conditions and camera poses. Even with recent advances in material acquisition allowing single or few image(s) capture [8, 9, 30, 42, 17, 55, 47], strong restrictions on the capture conditions are imposed. These methods typically require a flash light co-located with the camera as the only light source, and/or a fronto-parallel view of a flat sample. Even methods designed for capture using non-planar photographs [28] cannot handle significant texture distortion in the input photographs. Many recent material generation methods are trained from scratch on synthetic materials, limiting the generation diversity due to limited datasets [47, 1], as compared to general-purpose text-to-image diffusion models [31, 38, 39].
We propose to tackle these challenges with two new ideas. First, we create a dataset which contains 800K crops of synthetic scene renderings, textured with randomly assigned materials, with each crop paired with its ground truth material. Using this data, we can train our model for the “material picking task”, with various observation angles and distortion. We additionally use a text-to-material dataset [47] containing 800K pairs of text descriptions and associated ground truth material maps, encouraging material generation diversity and resulting in a multi-modal generator that can accept images, text or both.
Second, we re-purpose a text-to-video generation model to generate material maps instead. We use a Diffusion-Transformer (DiT) based architecture, which has been shown to be effective for high-quality video synthesis [6]. However, our target domain is materials, which we represent as a set of 2D maps (albedo, normal, height, roughness, metallicity). To adapt our base DiT model, trained on videos, to materials, we finetune it by considering each material map as a “frame” in a video sequence. This approach preserves the strong prior information in the video model, improving our method’s generalization and robustness.
We evaluate our model on both real and synthetic input images and compare it against the state-of-the-art methods for texture rectification [19], material picking [28] and text-to-material generation [47, 48]. We show that our approach generates materials that follow the input text prompt and/or match the appearance of the material sample in the input image, while correcting its distortions. In summary, we make the following contributions:
@@
  • We design a large-scale dataset of crops of material samples paired with the corresponding ground truth material maps, enabling our model to handle a range of viewing angles and distortion.
@@
2 Related Work
2.1 Material Acquisition and Generation
@@
While material acquisition has been a long standing challenge [15], lightweight material acquisition and generation have seen significant progress using machine learning. Various methods were proposed to infer PBR material maps from only a single or few photographs [8, 9, 10, 42, 17, 55]. However, these methods rely on specific capture condition using a flash light co-located with the camera location. Martin et al. [30] propose to use natural illumination but doesn’t support direct metallic and roughness map estimations. Further, these methods rely on the camera being fronto-parallel or very close to it. This kind of photographs require specific captures, making the use of in the wild photos for material creation challenging.
@@
As an alternative to create materials, generative model for materials were proposed. GAN-based approaches [18, 56] show that unconditional generation of materials is possible and can be used for material acquisition via optimization of their noise and latent spaces. Recent progress in generative model, and more specifically diffusion models [39], enabled more stable, diffusion based, material generators [48, 50]. Such diffusion models can also be used to support material acquisition tasks [47], for example when paired with ControlNet [54]. All these diffusion-based approaches either attempt to train the model from scratch, using solely synthetic material data [48, 47] or significantly alter the architecture of the original text-to-image model [50], preventing the use of the pre-existing priors in large scale image generation models [39], limiting their generalization and diversity. Further, image prompts are limited to fronto-parallel photographs, which requires a specific capture.
@@
Other methods leveraged transformers as a model for material generation [16, 25] but focused on procedural material, which relies on generating functional graph generation, a very different modality. These procedural representations have resolution and editability benefits, but cannot easily model materials with complex texture patterns in the wild. In contrast, our model supports generating materials from any image or text prompt and produces varied, high-quality material samples.
2.2 Material Extraction and Rectification
@@
Different methods were proposed to rectify textures or generally enable non-fronto-parallel textures as input. Some approaches [52, 51] aim to evaluate the materials in an image through a retrieval and optimization method. Given an image, they retrieve the geometries and procedural materials in databases to optimize their position and appearance via differentiable rendering [43, 53]. Closest to our work is Material palette [28], targeting material extraction from a single photo, not restricted to fronto parallel images. The method leverages Dreambooth [40] optimized through a LoRA  [24] on Stable Diffusion [39] to learn a “concept” for each material. This lets them generate a texture with a similar appearance to the target material and use a separate material estimation network to decompose the texture into material maps. However, this LoRA optimization step takes up to 3 minutes for each image, and we find that our approach reproduces better the target appearance.
@@
A related field is that of texture synthesis from real-world images. Wu et al. [49] present an automatic texture exemplar extraction based on Trimmed Texture CNN. VQGAN [12] achieves high resolution image-to-image synthesis with a transformer-based architecture. These methods however do not support the common occlusions and deformations that occur in natural images. To tackle this limitation, Hao et al. [19] propose to rectify occlusions and distortions in texture images via a conditional denoising UNet with an occlusion-aware latent transformer. We show that our approach yields better texture rectification and simultaneously generates material parameters.
2.3 Diffusion Models and Diffusion Transformers
@@
Diffusion models [23, 44, 45, 46] are state-of-the-art generative models, showing great results across various visual applications such as image synthesis and video generation. The core architecture of diffusion models progressed from simple UNets, incorporating self-attention and enhanced upscaling layers [11], prior-based text-to-image model [31, 38], a VAE [26] for latent diffusion models (LDM)  [39] and temporal attention layers for video generations [3, 2]. These image generation methods all rely on a U-net backbone, a convolutional-based encoder-decoder architecture.
Recently, transformer-based diffusion models, Diffusion Transformers (DiT) were proposed [35], benefiting from the scalability of Transformer models, removing the convolutions inductive bias. PixArt-$\alpha$ presents a DiT-based text-to-image that can synthesize high resolution images with low training cost. Stable Diffusion 3 [13] demonstrates that a multi-modal DiT model trained with Rectified Flow can achieve superior image synthesis quality. Compared to the U-net architecture, the DiT shows greater flexibility in the representation on the visual data, which is particularly important to video synthesis tasks. Sora [6], a DiT-based video diffusion model, encodes video sequences as tokens and uses transformers to denoise these visual tokens, demonstrating the ability to generate minute-long, high-resolution high-quality videos. We adapt a DiT-based video generation model for our purpose and show that it can be flexibly transformed into a multi-channel material generator.
3 Method
3.1 Diffusion Transformer
@@
To sample the original data distribution $p_{\text{data}}$ from the noise distribution, a reverse mapping $p(x_{0})=p(x_{T})\prod_{t=1}^{T}q(x_{t-1}|x_{t},\epsilon_{t})$ needs to be modeled where $\epsilon_{t}$ is the noise predicted at each step by a neural network $f_{\theta}$. The neural network $f_{\theta}$ is conditioned on the denoising step $t$ to predict the noise $\epsilon_{t}$, which is then used to reconstruct $x_{t-1}$ from $x_{t}$ in each reverse step [23]:
$\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1))}\left[\left\lVert\epsilon_{t}-f_{\theta}(x_{t};c,t)\right\rVert^{2}\right],$ (1)
@@
We use a Diffusion Transformer [35] architecture as a backbone to model $f_{\theta}$. The visual data $x\in\mathbb{R}^{T\times 3\times H\times W}$ is tokenized patch-wise, resulting in visual tokens $\hat{x}\in\mathbb{R}^{V\times D}$ where $H,W,T$ are the spatial and temporal dimensions of the video, $V$ is the number of tokens and $D$ is the feature dimension. Positional encoding is also added to $\hat{x}$ to specify spatial and temporal order. Any condition $c$ is also embedded as tokens $\hat{c}\in\mathbb{R}^{V^{\prime}\times D}$ where $V^{\prime}$ is the number of the tokens for conditional inputs. For example, when $c$ is a text, it is encoded by an pre-trained encoder [37] with additional embedding layers to map it into the same feature dimension $D$. The transformer $f_{\theta}(\hat{x}_{t};\hat{c},t)$ is trained to denoise each patch at timestep $t$. The final denoised patches $\hat{x}_{0}\in\mathbb{R}^{V\times D}$ are reassembled as visual data $x_{0}\in\mathbb{R}^{T\times 3\times H\times W}$ after decoding through linear layers. Since the number of tokens grows quickly with resolution, we use a variational autoencoder (VAE) model [35, 39] before the tokenizing process, producing a latent representation of $y\in\mathbb{R}^{T^{\prime}\times D^{\prime}\times H^{\prime}\times W^{\prime}}$ of the original data $x$ for the transformer to process.
3.2 Datasets
@@
For the Scenes dataset, we build a set of synthetic indoor scenes with planar floors, walls, and randomly placed 3D objects, such as cubes, spheres, cylinders, cones, and toruses, similar to random Cornell boxes [34]. Each object is randomly assigned a unique material from around 3,000 stationary (i.e., approximately shift invariant) materials. Using this approach we create a dataset of 100,000 high-resolution rendered images, with different kinds of light sources, including point lights and area lights, to simulate complex real-world illumination (see Fig. 2). We randomly place cameras to capture a wide variety of view points and maximize coverage.
@@
We further crop the rendered images to construct training data, including input images, corresponding material maps, binary material mask, and the material name as an optional text prompt. During cropping, we ensure that the dominant material occupies at least 70% of the region. Importantly, we rescale the material maps based on UV coordinates to ensure that the rendered crops and target material maps share a matching texture scale. After cropping, this dataset contains 800,000 text-image-mask-material tuples.
@@
As our Scenes dataset only contains stationary materials, it may fail to represent the full diversity of textures in the wild. To enhance the generalization capability, we use an additional Materials dataset [30], which we augment to 800,000 cropped material maps. We use the name of the materials as the text prompts for text-to-material generation. These data items can be thought of as text-material pairs. This additional data diversity leads to significant improvement for non-stationary textures in input photographs as discussed in Sec. 4.4.2.
3.3 Generative Material Model
We employ a pre-trained DiT-based text-to-video generative model, with an architecture similar to the one described in Brooks et al. [6], as a base model. We retarget it into a multi-channel material generator.
@@ 3.3 Generative Material Model
To retarget this model while preserving its learned prior knowledge, we stack the material maps $M$ (albedo map, normal map, height map, roughness map and metallicity map) into a “video” of 5 frames , and compute the temporal positional embedding assuming their time stamp interval is 1 e.g., fps=1. Since DiT flexibly generates tokenized data, as opposed to a U-net architecture [2], the number of frames it is able to produce is not fixed, allowing us to adapt the original video generator to generate the right number of “frames” to meet our requirement. Our proposed use of a video model is in contrast to image diffusion models which typically generate 3 channels (RGB) and need to be non-trivially adjusted to generate more channels [27].
To enable material generation from an image input, we consider the input image $I$ as the first frame, with the model generating the stacked material maps $M$ as the subsequent frames, similar to a video extension model. Recall that the input image can be captured with arbitrary camera pose, and may include perspective and distortions. Using this approach, the self-attention mechanism of the transformer ensures that the generated material parameters are aligned with each other, and allows for non-aligned pixels between the input image condition and the generated material maps. The model simply learns that all frames except the first need to be aligned. This property is key for texture rectification, which is challenging for convolution-based architectures, in which (approximate) pixel alignment between input and output is built in due to the convolution inductive bias.
@@
We additionally train our material generator to produce a segmentation mask for the dominant material in the crop. Typically, the user-provided crop is not entirely covered by a single material (see Fig. 4). Performing conservative cropping on an image may reduce the number of usable pixels, while using an additional segmentation mask requires additional user input or a separate segmentation model [41]. Instead, our model automatically identifies the dominant material [29] in the image. We add a mask $V$ to be inferred from the input image as the second frame. Our training data $x$ can thus be represented as $x=\text{stack}(I,V,M)$, where $x\in\mathbb{R}^{7\times 3\times H\times W}$; we have 7 RGB frames: input, mask, and five material maps. Noise $\epsilon_{t}$ is applied only to the last six frames occupied by $V$ and $M$, resulting in $x_{t}=\text{stack}(I,V_{t},M_{t})$, with the first frame (input image) remaining free of noise. Our objective from Eq. 1 is
@@
$\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1)}\left[\left\lVert\epsilon_{t}-f_{\theta}(x_{t};c,t)[-6:]\right\rVert^{2}\right],$ (2)
@@
where $c$ denotes the text input (material description). This process can be seen as a completion of the frames (mask and material channels) given the input image and text condition. The notation [6:][-6:][ - 6 : ] refers to the last 6 frames generated by the Transformer. When the input consists solely of $c$ without $I$, $x=\text{stack}(V,M)$ where $V$ is a uniformly white RGB image. The computation of the loss remains unchanged.
@@
(i) Text: “Wood, cherry wood laser cut bowling pin”
@@
(j) Text: “Marble granite, stylized light blue marble herringbone tiles”
@@
Figure 2: Our Scenes dataset. We build random scenes and render paired text/image-to-material dataset with 3K randomly sampled materials. In each row we show a 2K synthetic rendering, a crop with a dominant material, the material mask and corresponding material maps.
3.4 Training and Inference

### v3

# MaterialPicker: Multi-Modal DiT-Based Material Generation

@@ Abstract.
Material appearance, capture, generative models
*Corresponding authors: Xiaohe Ma (xiaohema1998@gmail.com).
Figure 1. We introduce MaterialPicker, a DiT-based model that generates high-quality materials, conditioned on image crops and/or text prompts. Our model accurately captures textures and material properties even from photographs of distorted or partially obscured surfaces. We demonstrate MaterialPicker by extracting material properties (albedo, normal, roughness, height and metallicity, shown in a column next to the input crops) from smartphone-captured photos, then applying these materials (as indicated by green arrows) in a 3D scene for photo-realistic rendering results.
1. Introduction
@@
Traditional material acquisition often requires tens or hundreds of photo samples under known light conditions and camera poses. Even with recent advances in material acquisition allowing single or few image(s) capture (Deschaintre et al., 2018, 2019; Martin et al., 2022; Shi et al., 2020a; Guo et al., 2021; Zhou and Kalantari, 2022; Vecchio et al., 2024a), restrictions on the capture conditions are imposed. These methods typically require a camera flash as the only light source, and/or a fronto-parallel view of a flat sample. Even methods designed for capture using non-planar photographs (Lopes et al., 2024) cannot handle significant texture distortion in the input photographs. Many recent material generation methods are trained from scratch on synthetic materials, limiting the generation diversity due to limited datasets (Vecchio et al., 2024a; Adobe, 2024), as compared to general-purpose text-to-image diffusion models (Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022).
We propose to tackle these challenges with two new ideas. First, we create a dataset which contains 800K crops of synthetic scene renderings, textured with randomly assigned materials, with each crop paired with its ground truth material. Using this data, we train our model for the “material picking" task. We additionally use a text-to-material dataset (Vecchio et al., 2024a) containing 800K pairs of text descriptions and associated ground truth material maps, encouraging material generation diversity and resulting in a multi-modal generator that can accept images, text or both.
Second, we re-purpose a text-to-video generation model to generate material maps instead. We use a Diffusion-Transformer (DiT) based architecture, which has been shown to be effective for high-quality video synthesis (Brooks et al., 2024). However, our target domain is materials, which we represent as a set of 2D maps (albedo, normal, roughness, height, metallicity). To adapt our base DiT model, trained on videos, to materials, we finetune it by considering each material map as a “frame” in a video sequence. This approach preserves the strong prior information in the video model, improving our method’s generalization and robustness.
We evaluate our model on both real and synthetic input images and compare it against the state-of-the-art methods for texture rectification (Hao et al., 2023), material picking (Lopes et al., 2024), material acquisition (Vecchio et al., 2024a) and text-to-material generation (Vecchio et al., 2024a, b). We show that our approach generates materials that follow the input text prompt and/or match the appearance of the material sample in the input image, while correcting its distortions. In summary, we make the following contributions:
@@
  • We design a large-scale dataset of crops of material samples paired with the corresponding ground truth material maps, enabling our model to handle a range of viewing angles and distortions.
@@
2. Related Work
2.1. Material Acquisition and Generation
@@
While material acquisition has been a long standing challenge (Guarnera et al., 2016), lightweight material acquisition and generation have seen significant progress using machine learning. Various methods were proposed to infer PBR material maps from only a single or few photographs (Deschaintre et al., 2018, 2019, 2020; Shi et al., 2020a; Guo et al., 2021; Zhou and Kalantari, 2022). However, these methods rely on specific capture condition using a flash light co-located with the camera location. Martin et al. (2022) propose to use natural illumination but doesn’t support direct metallic and roughness map estimations. Further, these methods rely on the camera being fronto-parallel or very close to it. This kind of photographs require specific captures, making the use of in the wild photos for material creation challenging.
@@
As an alternative to create materials, generative model for materials were proposed. GAN-based approaches (Guo et al., 2020; Zhou et al., 2022) show that unconditional generation of materials is possible and can be used for material acquisition via optimization of their noise and latent spaces. Recent progress in generative model, and more specifically diffusion models (Rombach et al., 2022), enabled more stable, diffusion based, material generators (Vecchio et al., 2024b; Xue et al., 2024). Such diffusion models can also be used to support material acquisition tasks (Vecchio et al., 2024a), for example when paired with ControlNet (Zhang et al., 2023). All these diffusion-based approaches either attempt to train the model from scratch, using solely synthetic material data (Vecchio et al., 2024b, a) or significantly alter the architecture of the original text-to-image model (Xue et al., 2024), preventing the use of the pre-existing priors in large scale image generation models (Rombach et al., 2022), limiting their generalization and diversity. Further, image prompts are limited to fronto-parallel photographs, which requires a specific capture.
@@
Other methods leveraged transformers as a model for material generation (Guerrero et al., 2022; Hu et al., 2023) but focused on procedural material, which relies on generating functional graph generation, a very different modality. These procedural representations have resolution and editability benefits, but cannot easily model materials with complex texture patterns in the wild. In contrast, our model supports generating materials from any image or text prompt and produces varied, high-quality material samples.
2.2. Material Extraction and Rectification
@@
Different methods were proposed to rectify textures or generally enable non-fronto-parallel textures as input. Some approaches (Yeh et al., 2022; Yan et al., 2023) aim to evaluate the materials in an image through a retrieval and optimization method. Given an image, they retrieve the geometries and procedural materials in databases to optimize their position and appearance via differentiable rendering (Shi et al., 2020b; Zhang et al., 2020). Closest to our work is Material Palette (Lopes et al., 2024), targeting material extraction from a single photo, not restricted to fronto parallel images. The method leverages Dreambooth (Ruiz et al., 2023) optimized through a LoRA  (Hu et al., 2021) on Stable Diffusion (Rombach et al., 2022) to learn a “concept” for each material. This lets them generate a texture with a similar appearance to the target material and use a separate material estimation network to decompose the texture into material maps. However, this LoRA optimization step takes up to 3 minutes for each image, and we find that our approach reproduces better the target appearance.
@@
A related field is that of texture synthesis from real-world images. Wu et al. (2020) present an automatic texture exemplar extraction based on Trimmed Texture CNN. VQGAN (Esser et al., 2021) achieves high resolution image-to-image synthesis with a transformer-based architecture. These methods however do not support the common occlusions and deformations that occur in natural images. To tackle this limitation, Hao et al. (2023) propose to rectify occlusions and distortions in texture images via a conditional denoising U-Net with an occlusion-aware latent transformer. We show that our approach yields better texture rectification and simultaneously generates material parameters.
2.3. Diffusion Models and Diffusion Transformers
@@
Diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Song et al., 2021) are state-of-the-art generative models, showing great results across various visual applications such as image synthesis and video generation. The core architecture of diffusion models progressed from simple U-Nets, incorporating self-attention and enhanced upscaling layers (Dhariwal and Nichol, 2021), prior-based text-to-image model (Nichol et al., 2021; Ramesh et al., 2022), a VAE (Kingma and Welling, 2022) for latent diffusion models (LDM)  (Rombach et al., 2022) and temporal attention layers for video generations (Blattmann et al., 2023b, a). These image generation methods all rely on a U-Net backbone, a convolutional-based encoder-decoder architecture.
[…]
3. Method
3.1. Diffusion Transformers
@@
To sample the original data distribution $p_{\text{data}}$ from the noise distribution, a reverse mapping $p(x_{0})=p(x_{T})\prod_{t=1}^{T}q(x_{t-1}|x_{t},\epsilon_{t})$ needs to be modeled where $\epsilon_{t}$ is the noise sampled at each step. A neural network $f_{\theta}$ is conditioned on the denoising step $t$ to predict the noise $\epsilon_{t}$, which is then used to reconstruct $x_{t-1}$ from $x_{t}$ in each reverse step (Ho et al., 2020):
(1) $\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,T))}\left[\left\lVert\epsilon_{t}-f_{\theta}(x_{t};c,t)\right\rVert^{2}\right],$
@@
We use a Diffusion Transformer (Peebles and Xie, 2023) architecture as a backbone to model $f_{\theta}$. The visual data $x\in\mathbb{R}^{F\times 3\times H\times W}$ is tokenized patch-wise, resulting in visual tokens $\hat{x}\in\mathbb{R}^{V\times D}$ where $H,W,F$ are the spatial and temporal dimensions of the video, $V$ is the number of tokens and $D$ is the feature dimension. Positional encoding is also added to $\hat{x}$ to specify spatial and temporal order. Any condition $c$ is also embedded as tokens $\hat{c}\in\mathbb{R}^{V^{\prime}\times D}$ where $V^{\prime}$ is the number of the tokens for conditional inputs. For example, when $c$ is a text, it is encoded by a pre-trained encoder (Radford et al., 2021) with additional embedding layers to map it into the same feature dimension $D$. The transformer $f_{\theta}(\hat{x}_{t};\hat{c},t)$ is trained to denoise each patch at timestep $t$. The final denoised patches $\hat{x}_{0}\in\mathbb{R}^{V\times D}$ are reassembled as visual data $x_{0}\in\mathbb{R}^{F\times 3\times H\times W}$ after decoding through linear layers. Since the number of tokens grows quickly with resolution, we use a variational autoencoder (VAE) model (Peebles and Xie, 2023; Rombach et al., 2022) before the tokenizing process, producing a latent representation of $y\in\mathbb{R}^{F^{\prime}\times D^{\prime}\times H^{\prime}\times W^{\prime}}$ of the original data $x$ for the transformer to process.
3.2. Datasets
@@
For the Scenes dataset, we build a set of synthetic indoor scenes with planar floors, walls, and randomly placed 3D objects, such as cubes, spheres, cylinders, cones, and toruses, similar to random Cornell boxes (Cohen and Greenberg, 1985). Each object is randomly assigned a unique material from around 3,000 stationary (i.e., approximately shift invariant) materials. We use the Blender implementation of the Disney Principled BSDF model (McAuley et al., 2012) for rendering the dataset and other visualizations in the paper, using base color (albedo), normal, roughness, metallic, and height maps, leaving other parameters as default. Using this approach we create a dataset of 100,000 high-resolution rendered images, with different kinds of light sources, including point lights and area lights, to simulate complex real-world illumination (see Fig. 2). We randomly place cameras to capture a wide variety of view points and maximize coverage.
@@
We further crop the rendered images to construct training data, including input images, corresponding material maps, binary material mask, and the material name as an optional text prompt. During cropping, we ensure that the dominant material occupies at least 70% of the region. Importantly, we rescale the material maps based on UV coordinates to ensure that the rendered crops and target material maps share a matching texture scale. After cropping, this dataset contains 800,000 text-image-mask-material tuples. We will share out dataset creation script, facilitating reproduction using public materials datasets (Vecchio and Deschaintre, 2024; Ma et al., 2023).
@@
As our Scenes dataset only contains stationary materials, it may fail to represent the full diversity of textures in the wild. To enhance the generalization capability, we use an additional Materials dataset (Martin et al., 2022), which we augment to 800,000 cropped material maps. We use the name of the materials as the text prompts for text-to-material generation. These data items can be thought of as text-material pairs. This additional data diversity leads to significant improvement for non-stationary textures in input photographs as discussed in Sec. 4.4.2.
3.3. Generative Material Model
We employ a pre-trained DiT-based text-to-video generative model as our base model, with an architecture similar to that of the publicly available HunyuanVideo model (Weijie et al., 2024). It follows a decoder-only Transformer structure with stacked self-attention blocks. The model takes both text and visual conditions as input: the input frames are first encoded by a 3DVAE encoder, producing a latent representation which is then corrupted by noise during training. Simultaneously, text prompts are processed by a T5-based encoder, producing text embeddings that are appended to the noisy latents. Additionally, timestep embeddings and spatial-temporal positional embeddings are added to the latent sequences to provide temporal and spatial context (i.e., each token’s frame number and position within the frame). The DiT backbone denoises the latent sequence, which is decoded into video frames using the 3DVAE decoder. We retarget this architecture into a multi-channel material generator.
@@ 3.3. Generative Material Model
To retarget the model while preserving its learned prior knowledge, we stack the material maps $M$ (albedo map, normal map, height map, roughness map and metallicity map) into a “video” of 5 frames, and compute the temporal positional embedding assuming their time stamp interval is 1 e.g., fps=1. Since DiT flexibly generates tokenized data, as opposed to a U-Net architecture (Blattmann et al., 2023a), the number of frames it is able to produce is not fixed, allowing us to adapt the original video generator to generate the right number of “frames” to meet our requirement. Note that we treat all material maps as “keyframes” (images) in the 3DVAE, with no motion prediction between the frames.
For image-conditioned material generation, we consider the input image $I$ as the first frame, with the model generating the stacked material maps $M$ as the subsequent frames, similar to a video extension model. This setup allows the transformer’s self-attention mechanism to jointly reason over both the input image and the predicted material maps, while tolerating pixel misalignment due to perspective distortion or varying camera poses.
Our decision to use a video-generation backbone is further motivated by its inherent capacity to enforce temporal consistency, which, in this context, translates to spatial alignment across material maps. The DiT model implicitly learns that all frames beyond the first should remain temporally coherent, a property that aligns well with the goal of generating consistent texture channels. Convolution-based architectures such as SDXL (Podell et al., 2023) led to worse results in our initial experiments; we hypothesize this is because they prefer (approximate) pixel alignment between input and output.
This design also avoids architectural modifications that disrupt pretrained knowledge. Image diffusion models are typically trained to generate 3 channels (RGB) and need to be non-trivially adjusted to generate more channels (Liu et al., 2023), or generate a single material map at a time and repurpose the input text prompt as a "switch"  (Zeng et al., 2024). Recent work (Vecchio et al., 2024a) modified the architecture to generate multiple maps, but had to train the model from random weights, missing the rich prior provided by large scale image datasets about material appearance. Our use of a video model enables the generation of multiple maps with minimally invasive architecture modifications, inheriting strong priors from pretrained video diffusion models.
Our solution preserves compatibility with inference-time techniques such as noise rolling (Vecchio et al., 2024a), TexSliders (Guerrero-Viu et al., 2024), SDEdit (Meng et al., 2021), etc., which could broaden its range of applications. Furthermore, since the DiT backbone operates on tokenized representations rather than fixed-size tensors (as in U-Net-based architectures), it is extensible, as additional maps (e.g., opacity maps) can be incorporated by appending new frames at the end of the token sequence. It also remains computationally efficient, since only a small number of frames is generated.
@@
We additionally train our material generator to produce a segmentation mask for the dominant material in the crop. Typically, the user-provided crop is not entirely covered by a single material (see Fig. 4). Performing conservative cropping on an image may reduce the number of usable pixels, while using an additional segmentation mask requires additional user input or a separate segmentation model (Sharma et al., 2023). Instead, our model automatically identifies the dominant material (Lu et al., 2009) in the image. We add a mask $S$ to be inferred from the input image as the second frame. Our training data $x$ can thus be represented as $x=\text{stack}(I,S,M)$, where $x\in\mathbb{R}^{7\times 3\times H\times W}$; we have 7 RGB frames: input, mask, and five material maps. Since mask, height, roughness, and metallic maps are single-channel, we convert them into RGB images before concatenating them with other frames. Noise $\epsilon_{t}$ is applied only to the last six frames occupied by $S$ and $M$, resulting in $x_{t}=\text{stack}(I,S_{t},M_{t})$, with the first frame (input image) remaining free of noise. Our objective from Eq. 1 is
@@
(2) $\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,T)}\left[\left\lVert\epsilon_{t}-f_{\theta}(x_{t};c,t)[-6:]\right\rVert^{2}\right],$
@@
where $c$ denotes the text material description. This process can be seen as frames completion (mask and material channels) given the input image and text condition. The notation [6:][-6:][ - 6 : ] refers to the last 6 frames generated by the Transformer. When the input consists solely of $c$ without $I$, $x=\text{stack}(S,M)$ where $S$ is a uniformly white RGB image. The computation of the loss remains unchanged.
@@
(i) Text: “Wood, cherry wood laser cut bowling pin"
@@
(j) Text: “Marble granite, stylized light blue marble herringbone tiles"
@@
Figure 2. Our Scenes dataset. We build random scenes and render paired text/image-to-material dataset with 3K randomly sampled materials. In each row we show a 2K synthetic rendering, a crop with a dominant material, the material mask and corresponding material maps.
3.4. Training and Inference